Calculate the molarity of the following solution.

0.300 mol of Na2S in 1.45 L of solution.

Calculate the molarity of the following solution.

0.350 mol of NaOH in 1.55 L of solution

Answer should be in M units.
Answer #1

molarity = no of moles/volume in liters

            = 0.3/1.45 = 0.2068M

molarity of Na2S = 0.2068M

molarity = no of moles/volume in liters

              = 0.35/1.55 = 0.2258M

molarity of NaOh is 0.2258M